OpenStudy (anonymous):

Two cars A and B are 200km apart and they are traveling toward each other at the speeds of 28km/h and 22 km/h respectively. When and where will they meet? 2. If the car A and B in problem 1 start from the same point, going toward the same direction, at what time will they be 18km apart? 3. If B starts at 1:00pm and gives chase starting at 1:15 from the same point as B, when and where will A overtake B? Thank you so much for answering!!

OpenStudy (phi):

3. If B starts at 1:00pm and gives chase starting at 1:15 from the same point as B, when and where will A overtake B? I think they left out A and it should read 3. If B starts at 1:00pm and A gives chase starting at 1:15 from the same point as B, when and where will A overtake B?

OpenStudy (phi):

So this question is more complicated. First notice that B starts at 1:00 pm and A does not start until 1:15 so B gets a "head start" can you figure out B's head start (how far did B get before A starts)

OpenStudy (anonymous):

If B starts at 1:00pm and A gives chase starting at 1:15 from the same point as B, when and where will A overtake B? ________ you're right

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

B gets a headstart of 15 mins??

OpenStudy (phi):

yes, now what is that in distance?

OpenStudy (anonymous):

22kph?

OpenStudy (phi):

what is the formula for distance?

OpenStudy (anonymous):

d= vt

OpenStudy (phi):

what is v and what is t?

OpenStudy (anonymous):

velocity - v time t

OpenStudy (phi):

yes, but for car B, in this problem, to figure out how far B got ahead of car A.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

how's it done?

OpenStudy (phi):

you need B's velocity and B's time (how long B drives before A starts)

OpenStudy (anonymous):

B drives 15mins before A

OpenStudy (anonymous):

then?

OpenStudy (phi):

you use d=v*t to figure out d (the distance) btw, v is km per hour so time has to be in hours (not minutes)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

how is it done?

OpenStudy (phi):

what is B's v? how long does B drive (in hours) before A starts ?

OpenStudy (anonymous):

22kph; 15mins before A

OpenStudy (phi):

what is 15 minutes in hours?

OpenStudy (anonymous):

1/4 hrs

OpenStudy (phi):

you know v and you know t so you can find d= v*t

OpenStudy (anonymous):

so if i multiply the two it wil result to d= 11/2

OpenStudy (phi):

now draw a picture showing where is A and where is B at 1:15

OpenStudy (anonymous):